# Malle’s conjecture and Brauer groups of stacks

### v2

@@ 6.1. Brauer group of $BG$
Note in particular that this lemma describes the Galois module structure of the finite group $\operatorname{Br}BG_{k^{\mathrm{sep}}}$. Moreover it allows us to identify the Picard scheme of $BG$ with $\widehat{G}$. Inspired by Sansuc [78, p. 42], we define $\operatorname{Br}_{e}BG=\{\beta\in\operatorname{Br}_{1}BG:\beta(e)=0\}$ where $e$ is the identity cocycle.
@@
Consider the fibre of the map $BG\to BG^{\text{ab}}$ over $e$. This is the $G^{\text{ab}}$-torsor $\pi:B[G,G]\to BG$. The type, as defined in [83, Def. 2.3.2], of this $G^{\text{ab}}$-torsor is a map $\operatorname{Hom}(G^{\text{ab}},\mathbb{G}_{\mathrm{m}})=\widehat{G}\to\operatorname{Pic}BG=\widehat{G}$. It is the identity map, which can be seen by unfolding the definitions.
@@
It follows from [83, Thm. 4.1.1] that $\cdot\cup[\pi]:\operatorname{H}^{1}(k,\widehat{G})\to\operatorname{Br}_{e}BG$ is an inverse to $r$ (this result is only stated for varieties, but the proof for stacks is analogous. To finish the lemma it suffices to notice that the map $BG(k)\to\operatorname{H}^{1}(k,G^{\text{ab}}),\chi\to[\pi](\chi)$ is by construction the map induced by $BG\to BG^{\text{ab}}$. ∎
@@ 6.2. Representation by central extensions
Recall that for an abelian group scheme $N$, the isomorphism classes of $N$-gerbes over a stack $\mathcal{X}$ are classified by the cohomology group $\operatorname{H}^{2}(\mathcal{X},N)$ [69, Thm. 12.2.8].
@@ 6.4. Marked central extensions
The Baer sum of two marked central extensions $1\to A\to E_{1}\to G\to 1$, $1\to A\to E_{2}\to G\to 1$ with respective markings $\mathcal{E}_{i}\subseteq E_{i}(-1)$ is the Baer sum extension [84, Tag 010I] $1\to A\to E_{1}+E_{2}\to G\to 1$. Note that $E_{1}+E_{2}$ is a subquotient of $E_{1}\times_{G}E_{2}$ and the marking $(\mathcal{E}_{1},\mathcal{E}_{2})\subseteq(E_{1}\times_{G}E_{2})(-1)$ defines a marking of $\mathcal{E}_{1}+\mathcal{E}_{2}\subseteq(E_{1}+E_{2})(-1)$.
@@ 6.4.1. The universal central marked extension
There are two main difference with our approach and the construction in [42, 94]. The first that we allow group schemes over fields, the second is that we since we define a marking as living in $E(-1)$ the universal marked central extension will be a non-constant group scheme even if $G$ is constant.
@@
The presence of $\zeta$ in the generators is to give the correct definition of the Galois action and to be able to define the morphism $\hat{U}(G,\mathcal{C})\to G:[\gamma]^{\zeta}\to\gamma(\zeta)$. This morphism preserves the relations and is equivariant with respect to the Galois action and is thus well-defined. It is a central extension by the same argument as in [94, Lem. 2.1].
@@
If $G$ is constant then $\hat{U}(G,\mathcal{C})(k^{\mathrm{sep}})$ agrees with what Wood [94, p. 7] denotes $\hat{U}_{k}(G,\mathcal{C})$ equipped with the Galois action given by $\sigma(u)=\chi_{\text{cycl}}(\sigma)\cdot u$, where $\sigma\in\Gamma_{k}$, $u\in\hat{U}_{k}(G,\mathcal{C})$, $\chi_{\text{cycl}}:\Gamma_{k}\to\widehat{\mathbb{Z}}^{\times}$ denotes the cyclotomic character and $\cdot$ is defined in loc. cit. page 7.
@@ 6.5. Orbifold Kummer exact sequence
An important consequence of this is that even though the Brauer group itself is not compatible with products, it turns out that the partially unramified Brauer group is, providing one is in a fair case.
@@ 6.7. Procedure for calculating $\operatorname{Br}_{\mathcal{C}}BG/\operatorname{Br}k$
If there are transcendental elements then by the orbifold Kummer sequence (Corollary 6.22) it suffices to calculate the orbifold cohomology $\operatorname{H}^{2,\mathrm{orb}}_{\mathcal{C}}(BG,\mathbb{Q}/\mathbb{Z}(1))$. One achieves this by calculating the universal marked central extension (Definition 6.15); if $G$ is constant then this is computable by [94, Thm. 2.5]. Then the orbifold cohomology is calculated using Proposition 6.18.
@@ 7.1. Ramification type
This has the following explicit description. An element $\varphi\in BG(k)$ is represented by a cocycle $\Gamma^{\text{tame}}_{k}\to G(k^{\mathrm{sep}})$. Moreover being continuous, it factors through some finite tamely ramified Galois extension $K/k$ of some ramification degree $e$. Enlarging $K$ if necessary, we assume that $K$ contains $\mu_{e}$. Let $\varpi$ be a uniformiser of $K/k$ and denote by $I_{K/k}\subseteq\operatorname{Gal}(K/k)$ the inertia group. By [84, Tag 09EE] the map
@@ 7.3.1. Via central extensions
be a tame central extension with corresponding Brauer group element $b_{E}\in\operatorname{Br}BG$ (see §6.2). Let $k\subseteq L$ be a field extension. Applying Galois cohomology, by [80, Prop. 5.7.43] we obtain the exact sequence of pointed sets
@@ 7.6.1. Cohomological argument
The existence of such a gerbe follows from the following. Poitou–Tate duality [67, Thm. 8.6.8] yields a perfect pairing
@@
for any number field $k$. It is well-known that the group (${}^{1}(k,\mu_{8})$, can be non-trivial for suitable $k$ [67, Thm. 9.1.3]; indeed, this is closely related to the Grunwald–Wang theorem already discussed. In such cases one deduces that (${}^{2}(k,\mathbb{Z}/8\mathbb{Z})$ is non-trivial. However $\operatorname{H}^{2}(k,\mathbb{Z}/8\mathbb{Z})$ classifies $\mathbb{Z}/8\mathbb{Z}$-gerbes, and a non-trivial element of (${}^{2}(k,\mathbb{Z}/8\mathbb{Z})$ exactly corresponds to a non-neutral $\mathbb{Z}/8\mathbb{Z}$-gerbe which is everywhere locally neutral, i.e. fails the Hasse principle. But then [70, Thm. 1.1] can be interpreted as the statement that the only obstruction to the Hasse principle for gerbes with constant abelian stabilisers is the Brauer–Manin obstruction, so this failure can be explained by the Brauer–Manin obstruction.
@@ 7.7. Brauer–Manin obstruction over global function fields
The map $\operatorname{Br}BG/\operatorname{Br}K\to\operatorname{Br}BG_{L}/\operatorname{Br}L$ need not be an isomorphism in general (even for $G=\mathbb{Z}/2\mathbb{Z}$). This gives further evidence for the naturality of the (partially) unramified Brauer group. Note that there is no reason to expect an analogue of Theorem 7.11 for non-constant group schemes in general. Non-constant $G$ are relevant in Malle’s conjecture when counting with unfair heights (see Conjecture 9.6). From our results one obtains the following significant strengthening of Theorem 7.4 for constant $G$.
@@ 8.1. Heights
We now prepare for considering Malle’s conjecture. We first define heights on $BG$. Our definition is inspired by the papers [32, 33], as well as the “$f$-discriminant” of Ellenberg and Venkatesh [40, §4.2] and Wood’s “counting function” [91, §2.1]. Let $G$ be a finite étale tame group scheme over a global field $k$.
@@
We call $H$ fair if $L$ is fair in the sense of Definition 3.30, i.e. if the minimal weight conjugacy classes $\mathcal{M}(H)$ generate $G$.
@@ 8.1.1. Examples
$\operatorname{ind}:G\to\mathbb{Z},\quad g\mapsto n-\#\{\text{number of orbits under multiplication by }g\}.$
The absolute value of the norm of the discriminant of the associated degree $n$ extension (see Lemma 2.15) is a height function corresponding to $\Delta$ (see [65, §7]).
@@ 8.2. Local Tamagawa measures
Let $H$ be a height associated to a big orbifold line bundle $L$. In [71, §2.2.1], Peyre defines a Tamagawa measure on $X(k_{v})$ for a Fano variety $X$ by taking an atlas of $X(k_{v})$ and gluing measures locally. He introduces a factor coming from a $v$-adic metric to ensure that these measures glue. Measures for other adjoint rigid line bundles are defined in [11, §3.3].
@@ 8.3. Mass formula
In this case $G$ has a unique finite smooth group scheme model $\mathcal{G}_{v}$ over $\mathcal{O}_{v}$; by abuse of notation we denote by $BG(\mathcal{O}_{v}):=B\mathcal{G}_{v}(\mathcal{O}_{v})$. If $G$ is constant, then $BG(\mathcal{O}_{v})$ is equivalent to the groupoid of homomorphisms $\Gamma_{\mathbb{F}_{v}}\to G$. Warning: $BG(\mathcal{O}_{v})\neq BG(k_{v})$ in general despite $BG$ being smooth and proper at $v$. This is because the valuative criterion for properness for stacks only guarantees a lift after a possibly ramified base-change; our replacement for this is Theorem 4.13.
@@
The equivalence easily follows from the isomorphism $\pi^{1}(\mathcal{O}_{v})\cong\operatorname{Gal}(\bar{\mathbb{F}}_{v}/\mathbb{F}_{v})$ The first equality follows from the fact that $H_{v}(\varphi_{v})=1$ for all $\varphi_{v}\in BG(\mathcal{O}_{v})$. The last equality is Lemma 8.8. ∎
@@
Peyre shows in [71, Lem. 2.2.1] that for a Fano variety $X$, we have $\tau_{v}(X(k_{v}))=\#X(\mathbb{F}_{v})/q_{v}^{\dim X}$ for all but finitely many places $v$. This is not true in our case; indeed we showed in Lemma 8.9 that #$\#BG(\mathbb{F}_{v})=1$. This difference all comes from the modified valuative criterion for properness for stacks.
@@ 8.5. Measure of the Brauer–Manin set
We now consider the Tamagawa measure of $BG(\mathbf{A}_{k})_{\mathcal{M}(L)}^{\operatorname{Br}}:=BG(\mathbf{A}_{k})_{\mathcal{M}(L)}^{\operatorname{Br}_{\mathcal{M}(L)}BG}$. To study this we assume that $H$ is fair, since this implies that $\operatorname{Br}_{\mathcal{M}(L)}BG/\operatorname{Br}k$ is finite (Corollary 6.30).
@@
$BG[\mathbf{A}_{k}]_{\mathcal{M}(L)}^{\mathscr{B}}\bigcap\left(\prod_{v\in S}BG[k_{v}]\prod_{v\notin S}BG[\mathcal{O}_{v}]_{\mathcal{C}}\right)=\left(\prod_{v\in S}BG[k_{v}]\right)^{\mathscr{B}}\prod_{v\notin S}BG[\mathcal{O}_{v}]_{\mathcal{C}}.$
The limit now follows from Definition 7.2 and absolute convergence. The non-vanishing is because $\left(\prod_{v\in S}BG[k_{v}]\right)^{\mathscr{B}}\prod_{v\notin S}BG[\mathcal{O}_{v}]_{\mathcal{C}}$ has positive measure by Corollary 8.11. ∎
@@
Lemma 8.21 leads to a natural hierarchy for $\tau_{H}(BG(\mathbf{A}_{k})_{\mathcal{M}(L)}^{\operatorname{Br}})$, with three cases, providing that $H$ is fair.
@@
    If $\operatorname{Br}_{\mathcal{M}(L)}BG=\operatorname{Br}_{\mathrm{un}}BG$, then $\tau_{H}(BG(\mathbf{A}_{k})_{\mathcal{M}(L)}^{\operatorname{Br}})$ is a sum of finitely many Euler products, where the Euler factors may differ at only finitely many places. This again follows from Theorem 5.23. This is the situation which has occurred so far in the literatue, for example in Wood’s paper [91] (see §10.7.4 for details).
@@ 9. Conjectures
where $\rho(X)=\operatorname{rank}\operatorname{Pic}X$, Peyre’s (renormalised) effective cone constant is denoted by $\alpha^{*}(X)$, and $\beta(X)=|\operatorname{H}^{1}(k,\operatorname{Pic}\bar{X})|=|\operatorname{Br}_{1}X/\operatorname{Br}k|$. (The constant $\beta(X)$ first appeared in [10, 11, 76].) Here $\tau$ is Peyre’s Tamagawa measure and $X(\mathbf{A}_{k})^{\operatorname{Br}}$ is the collection of adelic points orthogonal to $\operatorname{Br}X$. Peyre considered the anticanonical height only; for general line bundles we follow the dichotomy presented in [11, §3.4], as well as the predictions from [11, §3.3] (note that what Batyrev-Tschinkel call “$\mathcal{L}$-primitive” is now called “adjoint rigid” in the literature; in our paper we prefer the softer terminology “fair”).
Throughout this section, we let $H$ be a height function associated to a big orbifold line bundle $L=(\chi,w)$ on $BG$, for some non-trivial finite étale group scheme $G$ over a number field $k$ (see §8.1). We expect analogous conjectures to hold over global function fields providing $G$ is tame, but one needs to be slightly more careful with the statements due to the nature of asymptotic formulae over function fields, one often needs to take a weighted average over the count for different fixed height values (see e.g. [90, Thm. I.3] for example of this).
9.1. Fair heights
We say that $H$ (and $L$) is fair if the line bundle $L$ is adjoint rigid; equivalently if the collection $\mathcal{M}(L)$ of minimal weight elements generates $G$ (Definition 8.1). Recall that we denote by $BG[k]$ the set of isomorphism classes of elements of the groupoid $BG(k)$ (see §2.1 for basic properties of this groupoid).

### v3

@@ 6.1. Brauer group of $BG$
Note in particular that this lemma describes the Galois module structure of the finite group $\operatorname{Br}BG_{k^{\mathrm{sep}}}$. Moreover it allows us to identify the Picard scheme of $BG$ with $\widehat{G}$. Inspired by Sansuc [80, p. 42], we define $\operatorname{Br}_{e}BG=\{\beta\in\operatorname{Br}_{1}BG:\beta(e)=0\}$ where $e$ is the identity cocycle.
@@
Consider the fibre of the map $BG\to BG^{\text{ab}}$ over $e$. This is the $G^{\text{ab}}$-torsor $\pi:B[G,G]\to BG$. The type, as defined in [85, Def. 2.3.2], of this $G^{\text{ab}}$-torsor is a map $\operatorname{Hom}(G^{\text{ab}},\mathbb{G}_{\mathrm{m}})=\widehat{G}\to\operatorname{Pic}BG=\widehat{G}$. It is the identity map, which can be seen by unfolding the definitions.
@@
It follows from [85, Thm. 4.1.1] that $\cdot\cup[\pi]:\operatorname{H}^{1}(k,\widehat{G})\to\operatorname{Br}_{e}BG$ is an inverse to $r$ (this result is only stated for varieties, but the proof for stacks is analogous. To finish the lemma it suffices to notice that the map $BG(k)\to\operatorname{H}^{1}(k,G^{\text{ab}}),\chi\to[\pi](\chi)$ is by construction the map induced by $BG\to BG^{\text{ab}}$. ∎
@@ 6.2. Representation by central extensions
Recall that for an abelian group scheme $N$, the isomorphism classes of $N$-gerbes over a stack $\mathcal{X}$ are classified by the cohomology group $\operatorname{H}^{2}(\mathcal{X},N)$ [71, Thm. 12.2.8].
@@ 6.4. Marked central extensions
The Baer sum of two marked central extensions $1\to A\to E_{1}\to G\to 1$, $1\to A\to E_{2}\to G\to 1$ with respective markings $\mathcal{E}_{i}\subseteq E_{i}(-1)$ is the Baer sum extension [86, Tag 010I] $1\to A\to E_{1}+E_{2}\to G\to 1$. Note that $E_{1}+E_{2}$ is a subquotient of $E_{1}\times_{G}E_{2}$ and the marking $(\mathcal{E}_{1},\mathcal{E}_{2})\subseteq(E_{1}\times_{G}E_{2})(-1)$ defines a marking of $\mathcal{E}_{1}+\mathcal{E}_{2}\subseteq(E_{1}+E_{2})(-1)$.
@@ 6.4.1. The universal central marked extension
There are two main differences with our approach and the construction in [42, 96]. The first that we allow group schemes over fields, the second is that we since we define a marking as living in $E(-1)$ the universal marked central extension will be a non-constant group scheme even if $G$ is constant.
@@
The presence of $\zeta$ in the generators is to give the correct definition of the Galois action and to be able to define the morphism $\hat{U}(G,\mathcal{C})\to G:[\gamma]^{\zeta}\to\gamma(\zeta)$. This morphism preserves the relations and is equivariant with respect to the Galois action and is thus well-defined. It is a central extension by the same argument as in [96, Lem. 2.1].
@@
If $G$ is constant then $\hat{U}(G,\mathcal{C})(k^{\mathrm{sep}})$ agrees with what Wood [96, p. 7] denotes $\hat{U}_{k}(G,\mathcal{C})$ equipped with the Galois action given by $\sigma(u)=\chi_{\text{cycl}}(\sigma)\cdot u$, where $\sigma\in\Gamma_{k}$, $u\in\hat{U}_{k}(G,\mathcal{C})$, $\chi_{\text{cycl}}:\Gamma_{k}\to\widehat{\mathbb{Z}}^{\times}$ denotes the cyclotomic character and $\cdot$ is defined in loc. cit. page 7.
@@ 6.5. Orbifold Kummer exact sequence
An important consequence of this is that even though the Brauer group itself is not compatible with products, it turns out that the partially unramified Brauer group is, providing one is in a balanced case.
@@ 6.7. Procedure for calculating $\operatorname{Br}_{\mathcal{C}}BG/\operatorname{Br}k$
If there are transcendental elements then by the orbifold Kummer sequence (Corollary 6.22) it suffices to calculate the orbifold cohomology $\operatorname{H}^{2,\mathrm{orb}}_{\mathcal{C}}(BG,\mathbb{Q}/\mathbb{Z}(1))$. One achieves this by calculating the universal marked central extension (Definition 6.15); if $G$ is constant then this is computable by [96, Thm. 2.5]. Then the orbifold cohomology is calculated using Proposition 6.18.
@@ 7.1. Ramification type
This has the following explicit description. An element $\varphi\in BG(k)$ is represented by a cocycle $\Gamma^{\text{tame}}_{k}\to G(k^{\mathrm{sep}})$. Moreover being continuous, it factors through some finite tamely ramified Galois extension $K/k$ of some ramification degree $e$. Enlarging $K$ if necessary, we assume that $K$ contains $\mu_{e}$. Let $\varpi$ be a uniformiser of $K/k$ and denote by $I_{K/k}\subseteq\operatorname{Gal}(K/k)$ the inertia group. By [86, Tag 09EE] the map
@@ 7.3.1. Via central extensions
be a tame central extension with corresponding Brauer group element $b_{E}\in\operatorname{Br}BG$ (see §6.2). Let $k\subseteq L$ be a field extension. Applying Galois cohomology, by [82, Prop. 5.7.43] we obtain the exact sequence of pointed sets
@@ 7.6.1. Cohomological argument
The existence of such a gerbe follows from the following. Poitou–Tate duality [69, Thm. 8.6.8] yields a perfect pairing
@@
for any number field $k$. It is well-known that the group (${}^{1}(k,\mu_{8})$, can be non-trivial for suitable $k$ [69, Thm. 9.1.3]; indeed, this is closely related to the Grunwald–Wang theorem already discussed. In such cases one deduces that (${}^{2}(k,\mathbb{Z}/8\mathbb{Z})$ is non-trivial. However $\operatorname{H}^{2}(k,\mathbb{Z}/8\mathbb{Z})$ classifies $\mathbb{Z}/8\mathbb{Z}$-gerbes, and a non-trivial element of (${}^{2}(k,\mathbb{Z}/8\mathbb{Z})$ exactly corresponds to a non-neutral $\mathbb{Z}/8\mathbb{Z}$-gerbe which is everywhere locally neutral, i.e. fails the Hasse principle. But then [72, Thm. 1.1] can be interpreted as the statement that the only obstruction to the Hasse principle for gerbes with constant abelian stabilisers is the Brauer–Manin obstruction, so this failure can be explained by the Brauer–Manin obstruction.
@@ 7.7. Brauer–Manin obstruction over global function fields
The map $\operatorname{Br}BG/\operatorname{Br}K\to\operatorname{Br}BG_{L}/\operatorname{Br}L$ need not be an isomorphism in general (even for $G=\mathbb{Z}/2\mathbb{Z}$). This gives further evidence for the naturality of the (partially) unramified Brauer group. Note that there is no reason to expect an analogue of Theorem 7.11 for non-constant group schemes in general. Non-constant $G$ are relevant in Malle’s conjecture when counting with unbalanced heights (see Conjecture 9.6). From our results one obtains the following significant strengthening of Theorem 7.4 for constant $G$.
@@ 8.1. Heights
We now prepare for considering Malle’s conjecture. We first define heights on $BG$. Our definition is inspired by the papers [32, 33], as well as the “$f$-discriminant” of Ellenberg and Venkatesh [40, §4.2] and Wood’s “counting function” [93, §2.1]. Let $G$ be a finite étale tame group scheme over a global field $k$.
@@
We call $H$ balanced if $L$ is balanced in the sense of Definition 3.30, i.e. if the minimal weight conjugacy classes $\mathcal{M}(H)$ generate $G$.
@@ 8.1.1. Examples
$\operatorname{ind}:G\to\mathbb{Z},\quad g\mapsto n-\#\{\text{orbits under the action of }g\}.$
The absolute value of the norm of the relative discriminant of the associated degree $n$ extension (see Lemma 2.15) is a height function corresponding to $\Delta$ (see [66, §7]).
@@ 8.2. Local Tamagawa measures
Let $H$ be a height associated to a big orbifold line bundle $L$. In [73, §2.2.1], Peyre defines a Tamagawa measure on $X(k_{v})$ for a Fano variety $X$ by taking an atlas of $X(k_{v})$ and gluing measures locally. He introduces a factor coming from a $v$-adic metric to ensure that these measures glue. Measures for other adjoint rigid line bundles are defined in [11, §3.3].
@@ 8.3. Mass formula
In this case $G$ has a unique finite smooth group scheme model $\mathcal{G}_{v}$ over $\mathcal{O}_{v}$; by abuse of notation we denote by $BG(\mathcal{O}_{v}):=B\mathcal{G}_{v}(\mathcal{O}_{v})$. If $G$ is constant, then $BG(\mathcal{O}_{v})$ is equivalent to the groupoid of homomorphisms $\Gamma_{\mathbb{F}_{v}}\to G$. Warning: $BG(\mathcal{O}_{v})\neq BG(k_{v})$ in general despite $BG$ being smooth and proper at $v$. This is because the valuative criterion for properness for stacks [86, Tag 0CLK] only guarantees a lift after a possibly ramified base-change; our replacement for this is Theorem 4.13.
@@
The equivalence easily follows from the isomorphism $\pi_{1}(\mathcal{O}_{v})\cong\operatorname{Gal}(\bar{\mathbb{F}}_{v}/\mathbb{F}_{v})$ The first equality follows from the fact that $H_{v}(\varphi_{v})=1$ for all $\varphi_{v}\in BG(\mathcal{O}_{v})$. The last equality is Lemma 8.8. ∎
@@
Peyre shows in [73, Lem. 2.2.1] that for a Fano variety $X$, we have $\tau_{v}(X(k_{v}))=\#X(\mathbb{F}_{v})/q_{v}^{\dim X}$ for all but finitely many places $v$. This is not true in our case; indeed we showed in Lemma 8.9 that #$\#BG(\mathbb{F}_{v})=1$. This difference all comes from the modified valuative criterion for properness for stacks.
@@ 8.5. Measure of the Brauer–Manin set
We now consider the Tamagawa measure of $BG(\mathbf{A}_{k})_{\mathcal{M}(L)}^{\operatorname{Br}}:=BG(\mathbf{A}_{k})_{\mathcal{M}(L)}^{\operatorname{Br}_{\mathcal{M}(L)}BG}$. To study this we assume that $H$ is balanced, since this implies that $\operatorname{Br}_{\mathcal{M}(L)}BG/\operatorname{Br}k$ is finite (Corollary 6.30).
@@
$BG[\mathbf{A}_{k}]_{\mathcal{M}(L)}^{\mathscr{B}}\bigcap\left(\prod_{v\in S}BG[k_{v}]\prod_{v\notin S}BG[\mathcal{O}_{v}]_{\mathcal{M}(L)}\right)=\left(\prod_{v\in S}BG[k_{v}]\right)^{\mathscr{B}}\prod_{v\notin S}BG[\mathcal{O}_{v}]_{\mathcal{M}(L)}.$
The limit now follows from Definition 7.2 and absolute convergence. The non-vanishing is because $\left(\prod_{v\in S}BG[k_{v}]\right)^{\mathscr{B}}\prod_{v\notin S}BG[\mathcal{O}_{v}]_{\mathcal{M}(L)}$ has positive measure by Corollary 8.11. ∎
@@
Lemma 8.21 leads to a natural hierarchy for $\tau_{H}(BG(\mathbf{A}_{k})_{\mathcal{M}(L)}^{\operatorname{Br}})$, with three cases, providing that $H$ is balanced.
@@
    If $\operatorname{Br}_{\mathcal{M}(L)}BG=\operatorname{Br}_{\mathrm{un}}BG$, then $\tau_{H}(BG(\mathbf{A}_{k})_{\mathcal{M}(L)}^{\operatorname{Br}})$ is a sum of finitely many Euler products, where the Euler factors may differ at only finitely many places. This again follows from Theorem 5.23. This is the situation which has occurred so far in the literatue, for example in Wood’s paper [93] (see §LABEL:sec:balanced_wood for details).
@@ 9. Conjectures
where $\rho(X)=\operatorname{rank}\operatorname{Pic}X$, Peyre’s (renormalised) effective cone constant is denoted by $\alpha^{*}(X)$, and $\beta(X)=|\operatorname{H}^{1}(k,\operatorname{Pic}\bar{X})|=|\operatorname{Br}_{1}X/\operatorname{Br}k|$. (The constant $\beta(X)$ first appeared in [10, 11, 78].) Here $\tau$ is Peyre’s Tamagawa measure and $X(\mathbf{A}_{k})^{\operatorname{Br}}$ is the collection of adelic points orthogonal to $\operatorname{Br}X$. Peyre considered the anticanonical height only; for general line bundles we follow the dichotomy presented in [11, §3.4], as well as the predictions from [11, §3.3] (note that what Batyrev-Tschinkel call “$\mathcal{L}$-primitive” is now called “adjoint rigid” in the literature; in our paper we use the softer terminology “balanced”).
Throughout this section, we let $H$ be a height function associated to a big orbifold line bundle $L=(\chi,w)$ on $BG$, for some non-trivial finite étale group scheme $G$ over a number field $k$ (see §8.1). We expect analogous conjectures to hold over global function fields providing $G$ is tame, but one needs to be slightly more careful with the statements due to the nature of asymptotic formulae over function fields, one often needs to take a weighted average over the count for different fixed height values (see e.g. [92, Thm. I.3] for example of this).
9.1. Balanced heights
We say that $H$ (and $L$) is balanced if the line bundle $L$ is adjoint rigid; equivalently if the collection $\mathcal{M}(L)$ of minimal weight elements generates $G$ (Definition 8.1). Recall that we denote by $BG[k]$ the set of isomorphism classes of elements of the groupoid $BG(k)$ (see §2.1 for basic properties of this groupoid).